### v1

# Correlation between the first-reaction time and the acquired boundary local time

Abstract
We investigate the statistical correlation between the first-reaction time of a diffusing particle and its boundary local time accumulated until the reaction event. Since the reaction event occurs after multiple encounters of the particle with a partially reactive boundary, the boundary local time as a proxy for the number of such encounters is not independent of, but intrinsically linked to, the first-reaction time. We propose a universal theoretical framework to derive their joint probability density and, in particular, the correlation coefficient. To illustrate the dependence of these correlations on the boundary reactivity and shape, we obtain explicit analytical solutions for several basic domains. The analytical results are complemented by Monte Carlo simulations, which we employ to examine the role of interior obstacles on correlations in disordered media. Applications of these statistical results in chemical physics are discussed.
Keywords: first-reaction time, boundary local time, heterogeneous catalysis, diffusion-mediated reactions, Monte Carlo simulation
preprint: AIP/123-QED
I Introduction
The stochastic motion of particles and resulting encounters with their partners lie at the heart of numerous physical, chemical, and biological processes [1, 2, 3, 4, 5]. For diffusive particles inside a bounded domain $\Omega$ with Dirichlet boundary condition, the statistical characterization of these encounters through the distribution of first-passage times (FPTs) has been a central theme in statistical physics for decades [6, 7, 2, 8, 5, 9, 10, 3, 11, 13, 12, 14, 15, 4], where single or multiple targets can be either inside $\Omega$, or distributed on the boundary $\partial\Omega$}. In this classical framework, a target is considered perfectly absorbing, meaning the trajectory of a particle terminates upon the first encounter with the target. The FPT to such a target has been extensively analyzed, yielding profound insights into diffusion-limited kinetics [16, 17]. However, the first arrival of the particle on the reactive boundary does not necessarily achieve a successful absorption or reaction event. In most realistic scenarios, targets are not perfect sinks [4, 18, 19]. Examples range from trapping of reactants on catalytic surfaces [20, 21, 22] and binding of ligands to cellular receptors [23, 24] to the search of proteins for specific DNA sites [25, 26]. Microscopically, a molecule may interact with a reactive site multiple times due to insufficient overlapping between atom orbitals to form chemical bonds, thus being reflected back into the bulk after each non-reactive collision [27, 28]. Such reflected Brownian dynamics result in a random number of encounters with reactive sites before the successful absorption or reaction event occurs at the end. Therefore, the FPT can be expanded beyond the idealized scenario of a perfectly absorbing boundary to a more general concept of the first-reaction time (FRT) on a partially reactive boundary. In this setting, the propensity of the surface to react is characterized by a constant reactivity $\kappa$, while the ratio $q=\kappa/D$ quantifies the balance between diffusive and reactive fluxes on the boundary [21, 29]. The limit $q\to\infty$ recovers the perfect absorption such that the FRT is reduced to the FPT, while $q=0$ describes a perfectly reflecting, inert surface and the FRT turns out to be infinite.
Two particles with identical FRTs may have explored the partially reactive boundary to vastly different extents. A complete description of the reaction pathway thus requires a proper measure of this boundary exploration. This quantitative role is precisely fulfilled by the boundary local time (BLT) [30, 31, 32], which can be expressed as the rescaled number of encounters with the reactive boundary region $\Gamma$:
$\ell_{t}=\lim\limits_{\varepsilon\to 0}\varepsilon\,\mathcal{N}_{t}^{(\varepsilon)},$ (1)
where $\mathcal{N}_{t}^{(\varepsilon)}$ is the number of crossings of the thin layer $\Gamma_{\varepsilon}=\{\bm{x}\in\Omega:\left|\bm{x}-\Gamma\right|<\varepsilon\}$ of width $\varepsilon$ near the reactive region $\Gamma$ up to time $t$. Alternatively, the BLT $\ell_{t}$ quantifies the residence time in this boundary layer up to time $t$,
$\ell_{t}=\lim\limits_{\varepsilon\to 0}\frac{D}{\varepsilon}\int\limits_{0}^{t}\mathrm{d}t^{\prime}\,\Theta(\varepsilon-|\mathbf{X}_{t^{\prime}}-\Gamma|),$ (2)
where $D$ is the diffusion coefficient, $|\bm{x}-\Gamma|$ is the Euclidean distance between a point $\bm{x}$ and the reactive region $\Gamma$, and $\Theta(z)$ is the Heaviside step function: $\Theta(z)=1$ for $z>0$ and 0 otherwise. The BLT is used to construct the encounter-based approach to diffusion-controlled reactions [33], allowing one to consider different surface reaction mechanisms by imposing suitable stopping conditions [34, 35, 36, 37, 38, 39]. For instance, the reaction event on a partially reactive boundary occurs when $\ell_{t}$ exceeds an independent, exponentially distributed random threshold $\hat{\ell}$, satisfying $\mathbb{P}\{\hat{\ell}>\ell\}=e^{-q\ell}$. As a consequence, the first-reaction time on $\Gamma$ can be defined as:
$\tau=\inf\{t>0:\ell_{t}>\hat{\ell}\}$. (3)
This probabilistic formulation explicitly connects the macroscopic parameter $q$ to a microscopic stopping condition for the BLT. Still, the acquired BLT does not uniquely constrain the FRT. Even though two diffusive particles starting from the same location may acquire the same BLT, their FRTs can be substantially different. In other words, one cannot infer the FRT from the BLT, nor vice versa; in fact, these random variables provide complementary information on the reaction event. This raises a fundamental and yet unexplored question: How statistically correlated are the first-reaction time $\tau$ and the boundary local time $\ell_{\tau}$ acquired until the instance of reaction on $\Gamma$? In particular, the correlation coefficient $C_{q}=\mathrm{Corr}(\tau,\ell_{\tau})$ that measures the degree to which reaction time predicts boundary exploration (and vice versa), is crucial for deeper understanding of the diffusion-reaction pathways.
The paper is organized as follows. In Sec. II, we propose a general approach to derive the joint probability distribution of the FRT $\tau$ and the BLT $\ell_{\tau}$ and, in particular, their correlation coefficient. Section III describes Monte Carlo simulations, their validation for simple symmetric domains, and numerical results for disordered media with inert obstacles. We discuss further extensions and conclude the paper in Sec. IV.
II Theoretical results
Our analysis relies on the encounter-based approach [33]. We consider ordinary diffusion in a bounded domain $\Omega\subset\mathbb{R}^{d}$ whose smooth boundary ${\partial\Omega}=\Gamma\bigcup{\partial\Omega}_{N}$ is split into a partially reactive region $\Gamma$ and the remaining reflecting part ${\partial\Omega}_{N}$. For the boundary local time $\ell_{t}$ on $\Gamma$, we introduce the first-crossing time (FCT) $\mathcal{T}_{\ell}=\inf\{t>0:\ell_{t}>\ell\}$ of a fixed threshold $\ell$. The probability density $U(\ell,t|\bm{x}_{0})$ of the random variable $\mathcal{T}_{\ell}$ depends on the starting point $\bm{x}_{0}\in\Omega$, the threshold $\ell$, and the geometry of the confining domain $\Omega$. In particular, its Laplace transform
$\tilde{U}(\ell,p|\bm{x}_{0})=\left\langle e^{-p\mathcal{T}_{\ell}}\right\rangle=\int\limits_{0}^{\infty}\mathrm{d}t\,e^{-pt}U(\ell,t|\bm{x}_{0})\quad(p\geq 0),$ (4)
admits the spectral expansion [33] :
$\tilde{U}(\ell,p|\bm{x}_{0})=\sum_{k=0}^{\infty}e^{-\mu_{k}^{(p)}\ell}\ V_{k}^{(p)}(\bm{x}_{0})\int\limits_{\Gamma}\mathrm{d}\bm{x}\ V_{k}^{(p)}(\bm{x})\ ,$ (5)
where $\mu_{k}^{(p)}$ and $V_{k}^{(p)}(\bm{x})$ are the eigenvalues and eigenfunctions of the generalized Steklov problem:
$(p-D\Delta)V_{k}^{(p)}(\bm{x})$ $=0\qquad(\bm{x}\in\Omega),$ (6a)
$\partial_{n}V_{k}^{(p)}(\bm{x})$ $=\mu_{k}^{(p)}V_{k}^{(p)}\qquad(\bm{x}\in\Gamma),$ (6b)
$\partial_{n}V_{k}^{(p)}$ $=0\qquad(\bm{x}\in{\partial\Omega}_{N}),$ (6c)
where $\partial_{n}$ is the normal derivative oriented outward from the domain $\Omega$. The spectrum of this problem is known to be discrete [40], and the eigenpairs $\{\mu_{k}^{(p)},V_{k}^{(p)}\}$ are enumerated by $k=0,1,2,\ldots$ to order the eigenvalues $\mu_{k}^{(p)}$ into an increasing sequence. The moment of the FCT can be found from the moment-generating function $\tilde{U}(\ell,p|\bm{x}_{0})$:
$\langle\mathcal{T}^{m}_{\ell}\rangle=\int\limits_{0}^{\infty}\mathrm{d}t\ t^{m}\ U(\ell,t|\bm{x}_{0})=(-1)^{m}\left(\partial^{m}_{p}\tilde{U}(\ell,p|\bm{x}_{0})\right)_{p=0}.$ (7)
According to Eq. (3), the first-reaction time is the FCT of the random threshold $\hat{\ell}$, i.e., $\tau=\mathcal{T}_{\hat{\ell}}$. As a consequence, its probability density $H_{q}(t|\bm{x}_{0})$ can be obtained by averaging $U(\ell,t|\bm{x}_{0})$ over the exponential distribution of the threshold $\hat{\ell}$:
$H_{q}(t|\bm{x}_{0})=\langle U(\hat{\ell},t|\bm{x}_{0})\rangle=\int\limits_{0}^{\infty}\mathrm{d}\ell\ qe^{-q\ell}\ U(\ell,t|\bm{x}_{0}).$ (8)
Moreover, this formula indicates that
$\mathcal{P}_{q}(\ell,t|\bm{x}_{0})=qe^{-q\ell}\ U(\ell,t|\bm{x}_{0})\,$, (9)
is actually the joint probability density of $\ell_{\tau}$ and $\tau$. This conceptually important extension of $U(\ell,t|\bm{x}_{0})$ was unnoticed in previous works, and presents one of the main results of this paper. In particular, $H_{q}(t|\bm{x}_{0})$ is the marginal probability density of $\tau$, whereas $qe^{-q\ell}$ is the marginal probability density of $\ell_{\tau}=\hat{\ell}$.
For any integer $m>0$, one can express the $m$th moment of the first-reaction time as:
$\langle\tau^{m}\rangle$ $=\int\limits_{0}^{\infty}\mathrm{d}t\ t^{m}H_{q}(t|\bm{x}_{0})=(-1)^{m}\left(\partial_{p}^{m}\tilde{H}_{q}(p|\bm{x}_{0})\right)_{p=0}$
$=(-1)^{m}\int\limits_{0}^{\infty}\mathrm{d}\ell\ qe^{-q\ell}\left(\partial^{m}_{p}\tilde{U}(\ell,p|\bm{x}_{0})\right)_{p=0}$
$=\int\limits_{0}^{\infty}\mathrm{d}\ell\ qe^{-q\ell}\langle\mathcal{T}^{m}_{\ell}\rangle,$ (10)
where $\tilde{H}_{q}(p|\bm{x}_{0})$ is the Laplace transform of $H_{q}(t|\bm{x}_{0})$. An alternative calculation can be realized by the following recurrence relations for $m\geqslant 1$:
$D\Delta\langle\tau^{m}\rangle=-m\langle\tau^{m-1}\rangle$ $\qquad(\bm{x}\in\Omega),$ (11a)
$\partial_{n}\langle\tau^{m}\rangle+q\langle\tau^{m}\rangle=0$ $\qquad(\bm{x}\in\Gamma),$ (11b)
$\partial_{n}\langle\tau^{m}\rangle=0$ $\qquad(\bm{x}\in{\partial\Omega}_{N}),$ (11c)
with $\langle\tau^{0}\rangle=1$.
In turn, the $m$th moment of the acquired boundary local time $\ell_{\tau}$ is universal and independent of the domain $\Omega$:
$\langle\ell_{\tau}^{m}\rangle=\int\limits_{0}^{\infty}\mathrm{d}\ell\ \ell^{m}\int\limits_{0}^{\infty}\mathrm{d}t\ qe^{-q\ell}\ U(\ell,t|\bm{x}_{0})=\frac{m!}{q^{m}},$ (12)
where we used the normalization of the probability density $U(\ell,t|\bm{x}_{0})$, such that $\int_{0}^{\infty}\mathrm{d}t\ U(\ell,t|\bm{x}_{0})=1$. In particular, one gets the variance of $\ell_{\tau}$ as:
$\mathrm{Var}\{\ell_{\tau}\}=\frac{1}{q^{2}}.$ (13)
One also computes the crossing average as:
$\langle\tau\ell_{\tau}\rangle=\int\limits_{0}^{\infty}\mathrm{d}\ell\ \ell\int\limits_{0}^{\infty}\mathrm{d}t\ t\ qe^{-q\ell}\ U(\ell,t|\bm{x}_{0})=\int\limits_{0}^{\infty}\mathrm{d}\ell\ \ell\ qe^{-q\ell}\langle\mathcal{T}_{\ell}\rangle.$ (14)
Combining these results, we determine the correlation coefficient between $\tau$ and $\ell_{\tau}$:
$C_{q}=\mathrm{Corr}\{\tau,\ell_{\tau}\}=\frac{\langle\tau\ell_{\tau}\rangle-\langle\tau\rangle\langle\ell_{\tau}\rangle}{\sqrt{\mathrm{Var}\{\tau\}\mathrm{Var}\{\ell_{\tau}\}}}.$ (15)
Integrating Eq. (14) by parts, one can also obtain the following representation:
$\langle\tau\ell_{\tau}\rangle-\langle\tau\rangle\langle\ell_{\tau}\rangle=\int\limits_{0}^{\infty}\mathrm{d}\ell\ \ell e^{-q\ell}\partial_{\ell}\langle\mathcal{T}_{\ell}\rangle.$ (16)
As $\langle\mathcal{T}_{\ell}\rangle$ is an increasing function of $\ell$, the correlation is always positive.
II.1 Asymptotic behavior
Let us examine the asymptotic behavior of the correlation coefficient $C_{q}$.
In the limit of high reactivity ($q\to\infty$), both $\langle\tau\rangle$ and $\mathrm{Var}\{\tau\}$ approach their limits $\langle\tau_{\infty}\rangle$ and $\mathrm{Var}\{\tau_{\infty}\}$ corresponding to the first-passage time $\tau_{\infty}$ to a perfectly absorbing boundary region $\Gamma$. Moreover, the Laplace-transform relation
$\langle\tau\rangle=\int\limits_{0}^{\infty}\mathrm{d}\ell\,qe^{-q\ell}\,\langle\mathcal{T}_{\ell}\rangle$ (17)
(a particular case of Eq. (10) for $m=1$) reveals that the large-$q$ behavior of $\langle\tau\rangle$ corresponds to the small-$\ell$ behavior of $\langle\mathcal{T}_{\ell}\rangle$. As $\ell\to 0$, the mean first-crossing time $\langle\mathcal{T}_{\ell}\rangle$ also approaches $\langle\tau_{\infty}\rangle$: in fact, the first crossing of the threshold $\ell=0$ occurs at the first passage to the boundary. As a consequence, Eq. (14) implies that $\langle\tau\ell_{\tau}\rangle\approx\langle\tau_{\infty}\rangle/q$. Substituting these expressions into Eq. (15), we conclude that $C_{q}\to 0$ as $q\to\infty$. In other words, the first-reaction time $\tau$ and the acquired boundary local time $\ell_{\tau}$ become uncorrelated in this limit. This is not surprising: for a perfectly reactive boundary, one has $\ell_{\tau}=0$, whereas $\tau$ obeys the probability density $H_{\infty}(t|\bm{x}_{0})$, independently of $\ell_{\tau}$.
In the opposite limit of low reactivity ($q\to 0$), the mean FRT is dominated by slow reaction kinetics because the particle has to perform multiple reaction attempts and thus has enough time to diffuse through the bounded domain $\Omega$. Integrating Eq. (11a) with $m=1$ over $\Omega$, one gets
$|\Omega|=\int\limits_{\Omega}\mathrm{d}\bm{x}_{0}\,(-D\Delta\langle\tau\rangle)=qD\int\limits_{\Gamma}\mathrm{d}\bm{x}_{0}\,\langle\tau\rangle.$ (18)
As the starting point does not matter in the leading order, i.e., $\langle\tau\rangle\approx$ const, one has
$\langle\tau\rangle\approx\frac{|\Omega|}{qD|\Gamma|}\quad(q\to 0).$ (19)
In the same vein, one can integrate Eq. (11a) for the second moment ($m=2$) to get
$\int\limits_{\Omega}\mathrm{d}\bm{x}_{0}\,2\langle\tau\rangle=qD\int\limits_{\Gamma}\langle\tau^{2}\rangle,$ (20)
from which the same argument yields
$\langle\tau^{2}\rangle\approx 2\langle\tau\rangle\frac{|\Omega|}{qD|\Gamma|}\approx 2\biggl(\frac{|\Omega|}{qD|\Gamma|}\biggr)^{2}\quad(q\to 0),$ (21)
and thus $\sqrt{\mathrm{Var}\{\tau\}}\approx|\Omega|/(qD|\Gamma|)$. Finally, we employ again Eq. (17) to show that
$\langle\mathcal{T}_{\ell}\rangle\approx\frac{|\Omega|\ell}{D|\Gamma|}$ (22)
in the large-$\ell$ limit that corresponds to $q\to 0$. As a consequence, we get
$\langle\tau\ell_{\tau}\rangle\approx\frac{|\Omega|}{D|\Gamma|}\int\limits_{0}^{\infty}\mathrm{d}\ell\,\ell^{2}\,qe^{-q\ell}=\frac{2|\Omega|}{q^{2}D|\Gamma|}\,.$ (23)
Substituting these expressions into Eq. (15), we conclude that $C_{q}\to 1$ as $q\to 0$. This result is also not surprising: in the reaction-limited regime, the particle undergoes numerous reaction attempts, and the first-reaction time can be seen as the sum of random durations of diffusive motions in the bulk between successive encounters with the boundary; this sum is asymptotically proportional to the number of attempts, i.e., $\tau\sim\ell_{\tau}$, so that the two variables become fully correlated.
II.2 Special case for the starting point
It is instructive to consider the special case when the starting point is uniformly distributed over the reactive boundary region $\Gamma$. Substituting Eq. (17) into Eq. (18), we get
$\frac{|\Omega|}{q^{2}D|\Gamma|}=\int\limits_{0}^{\infty}\mathrm{d}\ell\,e^{-q\ell}\,\overline{\langle\mathcal{T}_{\ell}\rangle},$ (24)
where
$\overline{\langle\mathcal{T}_{\ell}\rangle}=\frac{1}{|\Gamma|}\int\limits_{\Gamma}\mathrm{d}\bm{x}_{0}\,\langle\mathcal{T}_{\ell}\rangle$ (25)
is the average (denoted by overline) of the mean FCT over the starting point uniformly distributed on $\Gamma$. Inverting the Laplace transform (24), we find
$\overline{\langle\mathcal{T}_{\ell}\rangle}=\frac{|\Omega|\ell}{D|\Gamma|}\,$, (26)
which is exact and valid for any confining domain. As a consequence, we obtain from Eq. (16):
$\frac{1}{|\Gamma|}\int\limits_{\Gamma}\mathrm{d}\bm{x}_{0}\bigl[\langle\tau\ell_{\tau}\rangle-\langle\tau\rangle\langle\ell_{\tau}\rangle\bigr]=$ $\int\limits_{0}^{\infty}\mathrm{d}\ell\,\,\ell\,\,e^{-q\ell}\,\,\partial_{\ell}\biggl(\frac{|\Omega|\ell}{D|\Gamma|}\biggr)$
$=$ $\frac{|\Omega|}{q^{2}D|\Gamma|},$ (27)
and the correlation coefficient in this case becomes:
$\overline{C_{q}}=\frac{|\Omega|}{qD|\Gamma|\sqrt{\overline{\mathrm{Var}\{\tau\}}}},$ (28)
where $\overline{\mathrm{Var}\{\tau\}}=\overline{\langle\tau^{2}\rangle}-\left(\overline{\langle\tau\rangle}\right)^{2}$ is the variance of the FRT with the uniform starting point on $\Gamma$. The boundary-averaged mean $\overline{\langle\tau\rangle}$ is given by Eq. (18):
$\overline{\langle\tau\rangle}=\frac{1}{|\Gamma|}\int\limits_{\Gamma}\mathrm{d}\bm{x}_{0}\langle\tau\rangle=\frac{|\Omega|}{qD|\Gamma|},$ (29)
whereas the second moment can be expressed from Eq. (20):
$\overline{\langle\tau^{2}\rangle}=\frac{1}{|\Gamma|}\int\limits_{\Gamma}\mathrm{d}\bm{x}_{0}\langle\tau^{2}\rangle=\frac{2|\Omega|}{qD|\Gamma|}{T}_{q},$ (30)
where ${T}_{q}=\frac{1}{|\Omega|}\int\limits_{\Omega}\mathrm{d}\bm{x}_{0}\langle\tau\rangle$. We get thus
$\overline{C_{q}}=\frac{1}{\sqrt{2qD{T}_{q}|\Gamma|/|\Omega|-1}},$ (31)
i.e., this correlation coefficient is expressed in terms of the volume-averaged mean FRT ${T}_{q}$. This exact relation exemplifies the transition between two asymptotic limits discussed in Sec. II.1: as $q\to 0$, one can use again Eq. (19), written as $T_{q}\simeq|\Omega|/(qD|\Gamma|)$, to see that $\overline{C_{q}}\to 1$; in turn, as $q\to\infty$, $T_{q}$ approaches the volume-averaged mean first-passage time $T_{\infty}$, and $\overline{C_{q}}$ vanishes. We will show below that the predicted decay $\overline{C_{q}}\propto q^{-1/2}$ is specific to the choice of the starting point on the reactive region $\Gamma$.
II.3 Example of a ball
To illustrate the above results, we consider diffusion in a ball $\Omega=\{\bm{x}\in\mathbb{R}^{3}:|\bm{x}|<R\}$ of radius $R$, with a partially reactive boundary ${\partial\Omega}=\Gamma$ (i.e., ${\partial\Omega}_{N}=\emptyset$), for which the Steklov eigenvalues and eigenfunctions are known explicitly (see, e.g., [33, 41]). The rotational symmetry implies that only the rotationally invariant eigenfunction $V_{0}^{(p)}(\bm{x})=\frac{1}{\sqrt{|\Gamma|}}\frac{i_{0}(\alpha|\bm{x}|)}{i_{0}(\alpha R)}$ contributes to Eq. (5), yielding
$\tilde{U}(\ell,p|\bm{x}_{0})$ $=e^{-\mu_{0}^{(p)}\ell}V_{0}^{(p)}(\bm{x}_{0})\int\limits_{\Gamma}\mathrm{d}\bm{x}\ V_{0}^{(p)}(\bm{x})$
$=e^{-\mu_{0}^{(p)}\ell}\frac{i_{0}(\alpha|\bm{x}_{0}|)}{i_{0}(\alpha R)},$ (32)
where $\alpha=\sqrt{p/D}$, $\mu_{0}^{(p)}=\alpha i_{0}^{\prime}(\alpha R)/i_{0}(\alpha R)$, $i_{0}(z)=\sinh(z)/z$, and $i_{0}^{\prime}(z)=\mathrm{d}i_{0}(z)/\mathrm{d}z$. For a starting point $|\bm{x}_{0}|=r$ inside the ball, one retrieves the Laplace transform of the FRT probability density:
$\tilde{H}_{q}(p|r)$ $=\int\limits_{0}^{\infty}\mathrm{d}\ell\ qe^{-q\ell}\ \tilde{U}(\ell,p|\bm{x}_{0})$
$=\frac{qi_{0}(\alpha r)}{\alpha i_{0}^{\prime}(\alpha R)+qi_{0}(\alpha R)},$ (33)
from which one gets
$\left\langle\tau\right\rangle(r)=$ $\frac{R^{2}-r^{2}}{6D}+\frac{R}{3Dq},$ (34a)
$\left\langle\tau^{2}\right\rangle(r)=$ $\frac{3r^{4}-10r^{2}R^{2}+7R^{4}}{180D^{2}}$
$+\frac{R\left(7R^{2}-5r^{2}\right)}{45D^{2}q}+\frac{2R^{2}}{9D^{2}q^{2}}$ (34b)
$\mathrm{Var}\{\tau\}(r)=$ $\frac{R^{4}-r^{4}}{90D^{2}}+\frac{10+4Rq}{90D^{2}q^{2}}R^{2}.$ (34c)
We have thus
$\left\langle\mathcal{T}_{\ell}\right\rangle(r)=\frac{R^{2}-r^{2}}{6D}+\frac{R\ell}{3D},$ (35)
while Eq. (14) implies
$\left\langle\tau\ell_{\tau}\right\rangle(r)$ $=\int\limits_{0}^{\infty}\mathrm{d}\ell\ \ell\ qe^{-q\ell}\left(\frac{R^{2}-r^{2}}{6D}+\frac{R\ell}{3D}\right)$
$=\frac{R^{2}-r^{2}}{6Dq}+\frac{2R}{3Dq^{2}},$ (36)
and thus
$\left\langle\tau\ell_{\tau}\right\rangle-\left\langle\tau\right\rangle\left\langle\ell_{\tau}\right\rangle=\frac{R}{3Dq^{2}}.$ (37)
Combining these results with Eq. (13), we determine the correlation coefficient between $\tau$ and $\ell_{\tau}$ from Eq. (15):
$C_{q}(r)=\frac{1}{\sqrt{1+2qR/5+(qR)^{2}(1-(r/R)^{4})/10}}.$ (38)
One can easily check the asymptotic behavior described in Sec. II.1: $C_{q}(r)\to 1$ as $q\to 0$ and $C_{q}(r)\to 0$ as $q\to\infty$. Moreover, we quantify now how fast or slow these limits are achieved; in particular, as $q\to\infty$, one gets $C_{q}(r)\propto q^{-1}$ for any $r<R$ and $C_{q}(R)\propto q^{-1/2}$, as we saw in Sec. II.4.
In Appendix A.1, we also derive the correlation coefficient for the starting point uniformly distributed inside the ball (we use the symbol $\circ$ to refer to this setting):
$C_{q}(\circ)=\frac{1}{\sqrt{1+2qR/5+13(qR)^{2}/175}}.$ (39)
One can see the same $2qR/5$ term in both Eqs. (38) and (39), implying the independence of $C_{q}$ on the starting point up to order $qR$ in the limit $qR\ll 1$. Similarly, the correlation coefficient for uniform starting points in the disk of radius $R$ is (see Appendix A.2)
$C_{q}(\circ)=\frac{1}{\sqrt{1+qR/2+5(qR)^{2}/48}}.$ (40)
II.4 Concentric spherical shell $(d\geq 2)$
Following the approach above, we now analyze diffusion inside a concentric spherical shell of radii $R$ and $L$. The inner boundary $\Gamma$ is partially reactive with the parameter $q>0$, while the outer boundary ${\partial\Omega}_{N}$ is inert. For this symmetric domain, the relation $\left\langle\tau\ell_{\tau}\right\rangle-\left\langle\tau\right\rangle\left\langle\ell_{\tau}\right\rangle=\frac{|\Omega|}{|\Gamma|Dq^{2}}$ holds. We present detailed calculations for the circular annulus ($d=2$) and the spherical shell ($d=3$) in Appendices A.3 and A.4, respectively.
It is instructive to inspect the correlation coefficient in the small-target limit $R\ll L$ and moderate reactivity. Expanding the correlation coefficient in powers of the small parameters $qR$ and $R/L$, we obtain in two dimensions (see Appendix A.3):
$C_{q}(r)=$ $1-qR\left(\log\left(\frac{L}{R}\right)-\frac{3}{4}\right)+\frac{(qR)^{2}}{32}\left[72\log\left(\frac{R}{L}\right)+48\log^{2}\left(\frac{R}{L}\right)\right$.
$\left.+8\log\left(\frac{r}{R}\right)\left(3+2\log\left(\frac{rR}{L^{2}}\right)\right)+27-16\left(\frac{r}{L}\right)^{2}+2\left(\frac{r}{L}\right)^{4}\right]+O\left(R^{3}\right),$ (41a)
$C_{q}(\circ)=$ $1-qR\left(\log\left(\frac{L}{R}\right)-\frac{3}{4}\right)+\frac{(qR)^{2}}{12}\left[6\log\left(\frac{L}{R}\right)\left(2\log\left(\frac{L}{R}\right)-3\right)+5\right]+O\left(R^{3}\right).$ (41b)
Curiously, the leading-order behavior of the correlation coefficient $C_{q}(r)$ is independent of the starting point $\bm{x}_{0}$ (note that $r=|\bm{x}_{0}|$ appears starting from the second-order term). Similarly, for diffusive particles near a tiny spherical target inside a large ball $(d=3)$, the initial position does not affect the correlation coefficient up to order $R^{3}$ (see Appendix A.4):
$C_{q}(r)=$ $1-qR+(qR)^{2}\left(1+\frac{9}{5qL}\right)-(qR)^{3}\left(1+\frac{18}{5qL}\right)$
$+\frac{qR^{4}}{100L^{6}r^{2}}\left(50L^{6}\left(2q^{3}r^{2}+q\right)+180L^{5}qr(3qr-1)+486L^{4}qr^{2}-100L^{3}r^{2}(qr+3)+5qr^{6}\right)+O\left(R^{5}\right),$ (42a)
$C_{q}(\circ)=$ $1-qR+(qR)^{2}\left(1+\frac{9}{5qL}\right)-(qR)^{3}\left(1+\frac{18}{5qL}\right)+\frac{qR^{4}(Lq(35Lq(5Lq+27)+459)-525)}{175L^{3}}+O\left(R^{5}\right).$ (42b)
These asymptotic results clearly show that the FRT and the BLT are strongly correlated in the small-target limit with moderate reactivity $q$ (such that $qR\ll 1$). We stress that the reactivity itself does not need to be small: one needs $q\ll q_{\mathrm{max}}$, with the maximal value $q_{\mathrm{max}}=1/R$ being large as $R\to 0$. While Eqs. (41, 42) were derived for a particular geometry of a spherical shell, we expect that the above conclusion remains valid in general.
III Numerical results
In this section, we employ Monte Carlo simulations to provide complementary insights into correlations between the FRT and the BLT. Sections III.1 and III.2 present the simulation algorithm and its validation, whereas Sections III.3, III.4, III.5 summarize our main numerical results for restricted diffusion in disordered media.
To investigate the role of geometric disorder, we consider domains containing interior inert obstacles (Fig. 1). We study three packing configurations: (a) a regular arrangement of identical disks of radius $R$ inside the unit disk, with the central disk being reactive and all the others being reflecting; (b) a random packing of identical non-overlapping disks of radius $R$ inside the unit disk; (c) a regular arrangement of identical balls of radius $R$ inside the unit ball, with the central ball being reactive.
III.1 Description of Monte Carlo simulations
To numerically investigate the joint statistics of the first-reaction time $\tau$ and the acquired boundary local time $\ell_{\tau}$, we perform extensive Monte Carlo simulations. For a given domain $\Omega$, starting points $\bm{x}_{0}$ are sampled from a uniform distribution inside $\Omega$. Each random trajectory $\mathbf{X}_{t}$ with $\mathbf{X}_{0}=\bm{x}_{0}$ is generated as follows. The walk-on-spheres (WOS) method [42] is employed for rapidly generating random points of a simulated trajectory inside $\Omega$, whereas the BLT remains unchanged. This sampling is repeated until the particle enters a thin boundary layer ${\partial\Omega}_{\varepsilon}=\{\bm{x}\in\Omega:|\bm{x}-{\partial\Omega}|<\varepsilon\}$ of a prescribed width $\varepsilon$. In this case, the escape-from-a-layer (EFL) approach [43] is used to accurately compute the increment of the BLT acquired until the escape of the particle from that layer. In this way, the most time-consuming simulation of reflected Brownian motion inside the layer is replaced by an equivalent escape event. Note that the BLT is incremented only in the vicinity of the reactive region $\Gamma$ whereas its increments near the reflecting part ${\partial\Omega}_{N}$ are ignored. After the escape from ${\partial\Omega}_{\varepsilon}$, the particle resumes its bulk diffusion simulated by WOS, and so on. The simulation of the trajectory is stopped when the acquired BLT $\ell_{t}$ exceeds an independently generated threshold $\hat{\ell}$, obeying the exponential distribution with a given parameter $q$. This stopping time is recorded as the FRT $\tau$, whereas the current value of the BLT is recorded as $\ell_{\tau}$. After that, the next trajectory is simulated from a newly generated starting point $\bm{x}_{0}$. Repeating this procedure $N$ times yields an empirical joint distribution of the FRT and the acquired BLT. From this dataset, we compute statistical moments (means, variances) and the correlation coefficient $C_{q}(\circ)$ according to Eq. (15). We generally use $N=10^{6}$ to ensure statistically robust estimates except for few cases explicitly specified. The proximity threshold $\varepsilon$ is chosen sufficiently small, $\varepsilon=10^{-3}$, to guarantee that the error in approximating the FRT and the BLT is negligible.
Figure 1: Computational domains with interior obstacles of the same radius $R$: the central one (in blue) is the reactive target, while the others are inert. (a) Regular square-lattice packing of $m\times m$ identical disks inside the unit disk, with $m=5$. The shown configuration has the maximal radius $R_{\mathrm{max}}$ such that all disks touch each other compactly without overlapping. (b) Random packing of 25 disks with the maximal radius $R_{\mathrm{max}}$ that corresponds to a fixed area occupation of 50%. The coordinates of all disks are provided in Appendix B. (c) Regular cubic-lattice packing of $m\times m\times m$ identical balls inside the unit ball with $m=3$, and the maximal radius $R_{\mathrm{max}}$.
III.2 Validation
Figure 2: Correlation coefficient $C_{q}(\circ)$ for (a) a unit ball and (b) a unit disk ($R=1$). Comparison between Monte Carlo results (MC, blue dots) and analytical expressions (39, 40) (orange curves). Numerical statistics are obtained from $N=10^{6}$ particles, with $D=1$. Green dashed lines indicate the asymptotic behaviors $C_{q}(\circ)\approx 5\sqrt{7/13}/(qR)$ for $d=3$ and $C_{q}(\circ)\approx 4\sqrt{3/5}/(qR)$ for $d=2$ in the limit $q\to\infty$.
Our Monte Carlo method is validated by comparing the correlation coefficients $C_{q}(\circ)$ obtained from simulations with the exact analytical expressions (39, 40) for uniform starting points inside the unit ball ($d=3$) or the unit disk ($d=2$). As shown in Fig. 2(a), the numerical results (blue dots) exhibit excellent agreement with Eq. (39) (orange curve) across a broad range of $q$, from $10^{-2}$ to $10^{3}$. In the limit $q\to 0$, one gets $C_{q}\to 1$, indicating that the FRT $\tau$ and the BLT $\ell_{\tau}$ become fully correlated. As $q$ increases, $C_{q}(\circ)$ decreases monotonically. In the large reactivity limit $q\to\infty$, a leading-order expansion of Eq. (39) yields the asymptotic behavior $C_{q}(\circ)\propto 1/q$. The same behavior is observed for the unit disk as shown in Fig. 2(b), where the simulation data match the exact solution given by Eq. (40).
Figure 3: Joint histograms of the FRT $\tau$ and the BLT $\ell_{\tau}$ for $q=10^{-2},10^{-1},10^{0},10^{1}$ obtained by Monte Carlo simulations with $N=10^{6}$ particles, where the starting points are uniform inside the unit ball. Both $\tau$ and $\ell_{\tau}$ are normalized by their respective mean values. In each panel, the scatter of blue dots represents individual Monte Carlo results of the pair $(\tau,\ell_{\tau})$, whose density approximated the underlying joint probability density $\mathcal{P}_{q}(\ell,t|\circ)$ from Eq. (9). The marginal distributions of $\tau/\langle\tau\rangle$ and $\ell_{\tau}/\langle\ell_{\tau}\rangle=q\ell_{\tau}$ are shown above and right of each joint plot.
Beyond the correlation coefficient itself, the joint histogram of $\tau$ and $\ell_{\tau}$ offers a more direct visual representation of their statistical coupling. Figure 3 presents such joint histograms for uniform starting points inside the unit ball at different reactivities. Both axes are normalized by their mean values given by Eq. (48a) and Eq. (12), respectively. The marginal distributions of $\tau$ and $\ell_{\tau}$ exhibit similar exponential decay across all $q$. For large $q$, the pairs $(\tau,\ell_{\tau})$ are widely scattered, reflecting weak correlations. In turn, for smaller reactivity, pairs concentrate along a straight line $\ell_{\tau}/\langle\ell_{\tau}\rangle\approx\tau/\langle\tau\rangle$, indicating the strong linear relation between $\tau$ and $\ell_{\tau}$ anticipated from the limit $C_{q}\to 1$ as $q\to 0$.
III.3 Regularly packed obstacles in 2d
Figure 4: Mean FRT and correlation coefficient for the domain shown in Fig. 1(a) with $m=5$. Numerical results are obtained by Monte Carlo simulations with $N=10^{6}$ particles that are initially uniformly distributed inside the domain. Symbols correspond to different reactivities: $q=10^{3}$ (squares), $q=10^{2}$ (downward triangles), $q=10^{1}$ (upward triangles), and $q=10^{0}$ (diamonds). (a) Mean FRT as a function of the scaled disk radius $R/R_{\mathrm{max}}$. The solid black curve represents the mean first-passage time (i.e., $q=\infty$), and colored dashed curves show the annulus approximation (44). (b) Correlation coefficient $C_{q}(\circ)$ as a function of scaled disk radius. Colored dashed curves correspond to the annulus approximation (66), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (43).
We employ Monte Carlo simulations to inspect the impact of geometric disorder onto the correlation coefficient. As illustrated by Fig. 1(a), we consider the unit disk containing $m^{2}$ identical disks arranged in a regular square-lattice packing. Given the non-overlapping condition, the maximal possible disk radius is $R_{\mathrm{max}}=\frac{L}{\sqrt{2}(m-1)+1}$, where $L=1$ is the outer boundary radius. Keeping the obstacle centers fixed, we vary their radii as $R=R_{\mathrm{max}}\times[0.01,0.1,0.2,0.3,0.5,0.7,0.9,0.99,1.0]$, to systematically investigate the disorder influence on both the first-reaction time and the correlation coefficient. The reactive target is the central disk, whereas the other disks are inert reflecting obstacles. To interpret the overall effect of the inert obstacles, we introduce an equivalent annulus domain, whose effective outer radius $L_{\mathrm{eff}}$ is determined by imposing the same area condition: $\pi(L_{\mathrm{eff}}^{2}-R^{2})=\pi(L^{2}-m^{2}R^{2})$, i.e.,
$L_{\mathrm{eff}}=\sqrt{L^{2}-(m^{2}-1)R^{2}}.$ (43)
Figure 4(a) compares the Monte Carlo results for mean FRTs obtained for $m=5$ with the annulus approximation (colored dashed lines) from Eq. (61b):
$\langle\tau\rangle(\circ)=\frac{L_{\mathrm{eff}}^{4}\log(L_{\mathrm{eff}}/R)}{2D(L_{\mathrm{eff}}^{2}-R^{2})}-\frac{3L_{\mathrm{eff}}^{2}-R^{2}}{8D}+\frac{L_{\mathrm{eff}}^{2}-R^{2}}{2qDR}.$ (44)
The black solid curve corresponds to the perfect absorption limit $q\to\infty$, i.e., the classical first-passage time. When the disk radius $R$ is small, the obstacles almost do not affect the search process, except for reducing the space available for diffusion, and the approximation agrees well with the simulations. Moreover, as the target (the central disk) has the same radius $R$, the mean FRT is large (we recall that it diverges as $1/(qR)$ as $R\to 0$). In turn, as $R$ increases, the mean FRT decreases, until reaching its minimum. The presence of this minimum is a peculiar feature of the considered arrangement of obstacles. In fact, in the limit case $R=R_{\mathrm{max}}$, as starting points are uniform in the domain, some of the particles cannot reach the reactive target, thus leading to infinite FRT. As this effect does not exist for the annulus domain, noticeable deviations appear at larger radii.
The correlation coefficient exhibits a similar trend. The equivalent annulus approximation from Eq. (66) remains accurate for $R/R_{\mathrm{max}}<0.3$ across all selected reactivities. In the high-reactivity limit $q\to\infty$, the power law $C_{q}(\circ)\propto 1/q$ is recovered, as seen from the separation between the blue ($q=10^{3}$) and orange ($q=10^{2}$) dashed curves shown in Fig. 4(b). As expected, when $q\to 0$, one has $C_{q}(\circ)\to 1$. As $R\to R_{\mathrm{max}}$, $C_{q}(\circ)$ shows a sharp decrease.
Figure 5: Role of the obstacle-lattice density $m$ at fixed reactivity $q=10^{2}$. Numerical results obtained by Monte Carlo simulations with $N=10^{6}$ particles that are initially uniformly distributed inside the domain shown in Fig. 1(a). Numerical results are shown for: $m=3$ (squares), $m=5$ (downward triangles), $m=7$ (upward triangles), and $m=9$ (diamonds). (a) Mean FRT as a function of the scaled disk radius $R/R_{\mathrm{max}}$. Colored dashed curves exhibit the annulus approximation (44) for each $m$. (b) Correlation coefficient $C_{q}(\circ)$ as a function of the scaled disk radius. Colored dashed curves exhibit the annulus approximation (66), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (43).
We also investigate the effect of obstacle density by changing $m$ value. Note that the maximal radius $R_{\mathrm{max}}$ itself depends on $m$ that also affects the horizontal axis in the figures, which is the scaled radius $R/R_{\mathrm{max}}$. As shown in Fig. 5(a), the annulus approximation works if $R/R_{\mathrm{max}}<0.3$, and then the mean FRT increases due to geometrical constraints. A larger density of obstacles (i.e., a larger $m$) results in a larger mean FRT. Although the area occupation fraction converges to a constant for a given ratio $R/R_{\mathrm{max}}$ as $m\to\infty$ (see Eq. (43)), the mean FRT diverges according to Eq. (44) since $R\lesssim R_{\mathrm{max}}\to 0$ as $m\to\infty$. Consistently, the correlation coefficient also increases with $m$ (Fig. 5(b)).
III.4 Randomly packed obstacles in two dimensions
Figure 6: Numerical results obtained by Monte Carlo simulations with $N=10^{5}$ particles that are initially uniformly distributed inside the domain shown in Fig. 1(b), with different reactivities: $q=10^{2}$ (squares), $q=10^{1}$ (downward triangles), $q=10^{0}$ (upward triangles), and $q=10^{-1}$ (diamonds). (a) Mean FPT as a function of the scaled disk radius $R/R_{\mathrm{max}}$. The black solid line represents the mean first-passage time (i.e., $q\to\infty$), and colored dashed lines exhibit the annulus approximation (44). (b) Correlation coefficient $C_{q}(\circ)$ as a function of the scaled disk radius. Colored dashed lines exhibit the annulus approximation (66), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (43).
In addition to the regular square-lattice packing, we examine another configuration in which the obstacles are packed randomly. To compare with previous results with $m=5$, we consider the configuration shown in Fig. 1(b) (see Appendix B for coordinates), with one reactive target at the center and 24 obstacles packed randomly. The maximal radius is determined by the 50% area occupation fraction. We consider the mean FRT (Fig. 6(a)) and the correlation coefficient (Fig. 6(b)) as before, where the annulus approximations (44) and (66) are also utilized with the effective radius given by Eq. (43). Compared with the regular packing, the annulus approximation also works when $R/R_{\mathrm{max}}<0.3$ for both mean FRT and correlation coefficient. The main difference arises in the limit $R\to R_{\mathrm{max}}$, in which the mean FRT is finite, and the correlation coefficient remains positive because the target remains accessible from any point in the domain (i.e., there is no geometric isolation of the target).
III.5 Regularly packed obstacles in three dimensions
Beyond two-dimensional configurations, we explore the effect of three-dimensional geometrical constraints via inert obstacles arranged on a regular cubic lattice inside the unit ball, as illustrated by Fig. 1(c). Again, given the non-overlapping condition, the maximal radius of each obstacle is $R_{\mathrm{max}}=L/(\sqrt{3}(m-1)+1)$, where $L=1$ is the outer boundary radius. Keeping the obstacle centers fixed, we vary their radii for numerical simulations. The overall effect of the obstacles is interpreted through an equivalent spherical shell domain, whose effective outer radius $L_{\mathrm{eff}}$ is determined by imposing the same volume condition, $\frac{4\pi}{3}(L_{\mathrm{eff}}^{3}-R^{3})=\frac{4\pi}{3}(L^{3}-m^{3}R^{3})$, i.e.,
$L_{\mathrm{eff}}=\sqrt[3]{L^{3}-(m^{3}-1)R^{3}}.$ (45)
Figure 7(a) compares the Monte Carlo results at $m=3$ for the mean FRT with this shell approximation (67b) (colored dashed lines). The black solid curve corresponds to the perfect absorption limit $q\to\infty$. Unlike the two-dimensional case, the approximation agrees well with simulations until a larger value of $R/R_{\mathrm{max}}$; especially for $q=10^{-1}$, there is still good agreement even at $R=R_{\mathrm{max}}$. In this limit, all particles can reach the reactive target within a finite FRT, since the three-dimension constraint does not isolate the target like in the two-dimensional case. The correlation coefficient (Fig. 7(b)) exhibits a similar trend. The equivalent shell approximation (72) remains accurate for $R/R_{\mathrm{max}}<0.3$. When the obstacle radius is close to the maximum, $C_{q}$ exhibits a decrease, but remains strictly positive.
Figure 7: Numerical results obtained by Monte Carlo simulations with $N=10^{4}$ particles that are initially uniformly distributed inside the domain shown in Fig. 1(c), with different reactivities: $q=10^{3}$ (squares), $q=10^{2}$ (downward triangles), $q=10^{1}$ (upward triangles), $10^{0}$ (diamonds), and $q=10^{-1}$ (dots). (a) Mean FRT as a function of the scaled ball radius $R/R_{\mathrm{max}}$. The black solid line represents the mean first-passage time (i.e., $q\to\infty$), and colored dashed lines exhibit the shell approximation (67b). (b) Correlation coefficient $C_{q}(\circ)$ as a function of the scaled ball radius. Colored dashed lines exhibit the shell approximation (72), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (45).
IV Discussion and conclusion
In this study, we have systematically investigated the statistical correlation between two fundamental quantities of diffusion-mediated surface reactions: the first-reaction time $\tau$ and the acquired boundary local time $\ell_{\tau}$. We developed a general theoretical framework for accessing their joint probability distribution and for calculating the correlation coefficient $C_{q}$. We derived explicit analytical expressions for $C_{q}$ in several basic domains, including the ball, the disk, the concentric spherical shell, and the concentric circular annulus. These exact results served as a benchmark for validating our Monte Carlo simulation technique based on the walk-on-spheres and the escape-from-a-layer algorithms. As these correlations are affected by the domain geometry, the analysis was then extended to disordered environments with interior obstacles (2D regular and random packings, as well as 3D regular packing). By employing concentric shell and annulus approximations, we examined the influence of geometric disorder by varying the number of obstacles and their radius. An interesting finding in two-dimensional regular configurations of obstacles is the emergence of a non-monotonic behavior in the mean FRT, a direct consequence of geometrical constraints. In the limiting geometric configurations that completely isolate a particle from the target, mean FRT diverges while the cross-moment vanishes. Across all other geometries, we confirmed the universal limiting behaviors of the correlation: $C_{q}\to 1$ as $q\to 0$, indicating perfect correlation when reactions are rare and boundary exploration is extensive; and $C_{q}\to 0$ as $q\to\infty$, reflecting the decoupling due to instantaneous absorption upon the first arrival. An important theoretical perspective is the analysis of the FRT, the BLT and their correlations for particles diffusing in porous media, in which pores are separated by narrow passages (as in the case of regularly arranged obstacles in two dimensions). Developing refined analytical tools for such poorly connected structures would be valuable.
This work opens several promising avenues for future research. For instance, the framework can be generalized to domains with multiple, disjoint reactive targets of different reactivities [44], where the joint distribution of multiple BLTs and their cross-correlations with the global FRT become relevant. Also, our findings on the strong dependence of $C_{q}$ on the geometry motivate a detailed study of the optimization problem: how does the correlation change with the shape, size, and spatial arrangement of reactive and inert regions? Can one achieve the desired correlation by re-arranging inert obstacles? Furthermore, extending the analysis to time-dependent reactivities $q(t)$ or spatially heterogeneous surface properties would bridge the studied model closer to applications in heterogeneous catalysis and dynamic biological interfaces. Exploring these directions will enhance our understanding of diffusion-mediated surface reactions and provide refined tools for designing and controlling processes in complex environments [45, 46] including porous media and cellular environments.
Acknowledgements
D.S.G. acknowledges the Alexander von Humboldt Foundation for support within a Bessel Prize award.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Calculations
For completeness, we provide intermediate results used in the calculation of the correlation coefficients for several basic domains, including the mean, second moment, and variance of the FRT, as well as the raw correlation between the FRT and the BLT. We recall that $(\circ)$ refers to the uniform starting point $\bm{x}_{0}\in\Omega$ with a constant density $\rho(\bm{x})=|\Omega|^{-1}$.
A.1 Ball
When the starting point is uniformly distributed inside a ball of radius $R$ (i.e., $\rho(\bm{x})=|\Omega|^{-1}=3/(4\pi R^{3})$), one has
$\tilde{U}(\ell,p|\circ)$ $=\frac{1}{|\Omega|}e^{-\mu_{0}^{(p)}\ell}\int\limits_{\Gamma}\mathrm{d}\bm{x}\ V_{0}^{(p)}(\bm{x})\int\limits_{\Omega}\mathrm{d}\bm{x}_{0}\ V_{0}^{(p)}(\bm{x}_{0})$
$=\frac{D\mu_{0}^{(p)}}{p|\Omega|}|\Gamma|e^{-\mu_{0}^{(p)}\ell},$ (46)
where $\Gamma={\partial\Omega}$, $\alpha=\sqrt{p/D}$, $\mu_{0}^{(p)}=\alpha i_{0}^{\prime}(\alpha R)/i_{0}(\alpha R)$, $V_{0}^{(p)}(\bm{x})=(1/\sqrt{|\Gamma|})i_{0}(\alpha|\bm{x}|)/i_{0}(\alpha R)$, and we used the identity $\int_{\Omega}\mathrm{d}\bm{x}_{0}\ V_{0}^{(p)}(\bm{x}_{0})=D\mu_{0}^{(p)}\int_{\partial\Omega}\mathrm{d}\bm{x}\ V_{0}^{(p)}(\bm{x})/p$. Also, one has
$\tilde{H}_{q}(p|\circ)$ $=\frac{1}{|\Omega|}\int\limits_{\Omega}\mathrm{d}\bm{x}_{0}\ \tilde{H}_{q}(p|\bm{x}_{0})=\int\limits_{0}^{\infty}\mathrm{d}\ell\,qe^{-q\ell}\tilde{U}(\ell,p|\circ)$
$=\frac{qD\mu_{0}^{(p)}|\Gamma|}{p|\Omega|(q+\mu_{0}^{(p)})},$ (47)
from which one gets the moments:
$\left\langle\tau\right\rangle(\circ)$ $=\frac{R^{2}}{15D}+\frac{R}{3Dq},$ (48a)
$\left\langle\tau^{2}\right\rangle(\circ)$ $=\frac{2R^{2}}{315D^{2}q^{2}}(35+14qR+2q^{2}R^{2}),$ (48b)
$\mathrm{Var}\{\tau\}(\circ)$ $=\frac{R^{2}}{1575D^{2}q^{2}}(13q^{2}R^{2}+70qR+175).$ (48c)
From Eq. (46), we have
$\left\langle\mathcal{T}_{\ell}\right\rangle(\circ)=\frac{R^{2}}{15D}+\frac{R\ell}{3D}.$ (49)
Thus the cross-moment reads
$\left\langle\tau\ell_{\tau}\right\rangle(\circ)$ $=\int\limits_{0}^{\infty}\mathrm{d}\ell\ \ell\ qe^{-q\ell}\left(\frac{R^{2}}{15D}+\frac{R\ell}{3D}\right)$
$=\frac{R^{2}}{15Dq}+\frac{2R}{3Dq^{2}},$ (50)
and we get
$\left\langle\tau\ell_{\tau}\right\rangle-\left\langle\tau\right\rangle\left\langle\ell_{\tau}\right\rangle=\frac{R}{3Dq^{2}},$ (51)
which is the same as Eq. (37) for a fixed starting point.
A.2 Disk
Consider diffusing particles inside the disk of radius $R$, whose boundary $\Gamma={\partial\Omega}$ is partially reactive with $q>0$. For the uniform starting point, both Eqs. (46) and (47) remain valid, with $\alpha=\sqrt{p/D}$, $\mu_{0}^{(p)}=\alpha I_{1}(\alpha R)/I_{0}(\alpha R)$, $V_{0}^{(p)}(\bm{x})=(1/\sqrt{|\Gamma|})I_{0}(\alpha|\bm{x}|)/I_{0}(\alpha R)$, where $I_{n}(z)$ is the modified Bessel function of the first kind. In turn, when the starting point is fixed, one has
$\tilde{U}(\ell,p|\bm{x}_{0})$ $=e^{-\mu_{0}^{(p)}\ell}V_{0}^{(p)}(\bm{x}_{0})\int\limits_{\Gamma}\mathrm{d}\bm{x}\ V_{0}^{(p)}(\bm{x})$
$=e^{-\mu_{0}^{(p)}\ell}\frac{I_{0}(\alpha|\bm{x}_{0}|)}{I_{0}(\alpha R)},$ (52)
and thus
$\tilde{H}_{q}(p|\bm{x}_{0})$ $=\int\limits_{0}^{\infty}\mathrm{d}\ell\ qe^{-q\ell}\ \tilde{U}(\ell,p|\bm{x}_{0})$
$=\frac{qI_{0}(\alpha|\bm{x}_{0}|)}{\alpha I_{1}(\alpha R)+qI_{0}(\alpha R)}.$ (53)
The mean FRTs are
$\langle\tau\rangle(r)$ $=\frac{R^{2}-r^{2}}{4D}+\frac{R}{2Dq},$ (54a)
$\langle\tau\rangle(\circ)$ $=\frac{R^{2}}{8D}+\frac{R}{2Dq}.$ (54b)
Solving Eqs. (11) for $m=2$ implies
$\langle\tau^{2}\rangle(r)$ $=\frac{R^{2}}{2D^{2}q^{2}}+\frac{3R^{3}-2r^{2}R}{8D^{2}q}+\frac{r^{4}-4r^{2}R^{2}+3R^{4}}{32D^{2}},$ (55a)
$\langle\tau^{2}\rangle(\circ)$ $=\frac{R^{2}}{2D^{2}q^{2}}+\frac{R^{3}}{4D^{2}q}+\frac{R^{4}}{24D^{2}},$ (55b)
and thus the variances read
$\mathrm{Var}\{\tau\}(r)$ $=\frac{q^{2}\left(R^{4}-r^{4}\right)+4qR^{3}+8R^{2}}{32D^{2}q^{2}},$ (56a)
$\mathrm{Var}\{\tau\}(\circ)$ $=\frac{R^{2}(qR(5qR+24)+48)}{192D^{2}q^{2}}.$ (56b)
Moreover, we have
$\langle\mathcal{T}_{\ell}\rangle(r)$ $=\frac{R^{2}-r^{2}}{4D}+\frac{R\ell}{2D},$ (57a)
$\langle\mathcal{T}_{\ell}\rangle(\circ)$ $=\frac{R^{2}}{8D}+\frac{R\ell}{2D}.$ (57b)
As a result, we get the cross-moments
$\langle\tau\ell_{\tau}\rangle(r)$ $=\int\limits_{0}^{\infty}\mathrm{d}\ell\,\ell\ qe^{-q\ell}\langle\mathcal{T}_{\ell}\rangle(r)=\frac{q\left(R^{2}-r^{2}\right)+4R}{4Dq^{2}},$ (58a)
$\langle\tau\ell_{\tau}\rangle(\circ)$ $=\int\limits_{0}^{\infty}\mathrm{d}\ell\,\ell\ qe^{-q\ell}\langle\mathcal{T}_{\ell}\rangle(\circ)=\frac{R(qR+8)}{8Dq^{2}},$ (58b)
so that
$\langle\tau\ell_{\tau}\rangle-\langle\tau\rangle\langle\ell_{\tau}\rangle=\frac{|\Omega|}{|\partial\Omega|Dq^{2}}=\frac{R}{2Dq^{2}}$ (59)
for both cases of a fixed and uniformly distributed starting point. The correlation coefficient $C_{q}$ then follows as
$C_{q}(r)$ $=\frac{1}{\sqrt{1+qR/2+(qR)^{2}\left(1-(r/R)^{4}\right)/8}},$ (60a)
$C_{q}(\circ)$ $=\frac{1}{\sqrt{1+qR/2+5(qR)^{2}/48}}.$ (60b)
A.3 Annulus
Consider particles diffusing inside two-dimensional concentric circular annulus of radii $R<L$, where the inner boundary $\Gamma$ (at $r=R$) is partially reactive with $q>0$ and the outer one ${\partial\Omega}_{N}$ (at $r=L$) is inert. Solving Eqs. (11) for $m=1$, we get the mean FPTs:
$\langle\tau\rangle(r)$ $=\frac{L^{2}}{2D}\log\left(\frac{r}{R}\right)+\frac{R^{2}-r^{2}}{4D}+\frac{L^{2}-R^{2}}{2qDR},$ (61a)
$\langle\tau\rangle(\circ)$ $=\frac{L^{4}\log(L/R)}{2D(L^{2}-R^{2})}-\frac{3L^{2}-R^{2}}{8D}+\frac{L^{2}-R^{2}}{2qDR}.$ (61b)
Solving Eqs. (11) for $m=2$ yields
$\langle\tau^{2}\rangle(r)$ $=\frac{1}{32D^{2}q^{2}R^{2}}\Bigg\{\begin{aligned} &4L^{4}(4-3qR)-8L^{2}R\left(qr^{2}(1-qR)+qR^{2}(qR-3)+4R\right)\\ &+R^{2}\left(q^{2}\left(r^{4}-4r^{2}R^{2}+3R^{4}\right)+4q\left(2r^{2}R-3R^{3}\right)+16R^{2}\right)\\ &+4qRL^{2}\left[4L^{2}\log\left(\frac{L}{R}\right)\left(qR\log\left(\frac{r}{R}\right)+1\right)+\log\left(\frac{r}{R}\right)\left(L^{2}(4-3qR)-2R\left(qr^{2}-qR^{2}+2R\right)\right)\right]\Bigg\},\end{aligned}$ (62a)
$\langle\tau^{2}\rangle(\circ)$ $=\frac{1}{48D^{2}q^{2}R^{2}\left(L^{2}-R^{2}\right)}\Bigg\{\begin{aligned} &\left(L^{2}-R^{2}\right)\left(L^{4}\left(17q^{2}R^{2}-36qR+24\right)-L^{2}R^{2}\left(13q^{2}R^{2}-48qR+48\right)+2R^{4}\left(q^{2}R^{2}-6qR+12\right)\right)\\ &-12L^{4}qR\log\left(\frac{L}{R}\right)\left(L^{2}(3qR-4)-2L^{2}qR\log\left(\frac{L}{R}\right)-2R^{2}(qR-2)\right)\Bigg\}.\end{aligned}$ (62b)
By replacing $1/q$ as $\ell$ in $\langle\tau\rangle$, one gets
$\langle\mathcal{T}_{\ell}\rangle(r)$ $=\frac{L^{2}}{2D}\log\left(\frac{r}{R}\right)+\frac{R^{2}-r^{2}}{4D}+\frac{L^{2}-R^{2}}{2DR}\ell,$ (63a)
$\langle\mathcal{T}_{\ell}\rangle(\circ)$ $=\frac{L^{4}\log(L/R)}{2D(L^{2}-R^{2})}-\frac{3L^{2}-R^{2}}{8D}+\frac{L^{2}-R^{2}}{2DR}\ell.$ (63b)
The cross-moment reads
$\langle\tau\ell_{\tau}\rangle(r)$ $=\frac{L^{2}-R^{2}}{Dq^{2}R}+\frac{R^{2}-r^{2}}{4Dq}+\frac{L^{2}\log(r/R)}{2Dq},$ (64a)
$\langle\tau\ell_{\tau}\rangle(\circ)$ $=\frac{L^{2}-R^{2}}{Dq^{2}R}+\frac{R^{2}-3L^{2}}{8Dq}+\frac{L^{2}\log(L/R)}{2Dq(L^{2}-R^{2})},$ (64b)
so that
$\langle\tau\ell_{\tau}\rangle-\langle\tau\rangle\langle\ell_{\tau}\rangle=\frac{|\Omega|}{|\partial\Omega|Dq^{2}}=\frac{L^{2}-R^{2}}{2Dq^{2}R}$ (65)
in both cases. Combining these results, one determines $C_{q}(\circ)$ from Eq. (15) for the annulus approximation used in Fig. 4(b), 5(b), and 6(b):
$C_{q}(\circ)$ $=\frac{4\sqrt{3}(L^{2}-R^{2})}{\sqrt{\begin{aligned} &L^{4}(qR(41qR-72)+48)-2L^{2}R^{2}(qR(17qR-48)+48)+R^{4}(qR(5qR-24)+48)\\ &+24L^{4}qR\log\left(\frac{L}{R}\right)\left(\frac{2L^{2}qR\left(L^{2}-2R^{2}\right)\log\left(\frac{L}{R}\right)}{\left(L^{2}-R^{2}\right)^{2}}-3qR+4\right)\end{aligned}}}\ .$ (66)
A.4 Spherical shell
For the three-dimensional concentric spherical shell of radii $R<L$ with the reactive inner boundary and the inert outer one, the mean FRTs read
$\langle\tau\rangle(r)$ $=\frac{L^{3}-R^{3}}{3DqR^{2}}+\frac{2L^{3}(r-R)-r^{3}R+rR^{3}}{6DrR},$ (67a)
$\langle\tau\rangle(\circ)$ $=\frac{L^{3}-R^{3}}{3DqR^{2}}+\frac{(L-R)^{2}\left(5L^{3}+6L^{2}R+3LR^{2}+R^{3}\right)}{15DR\left(L^{2}+LR+R^{2}\right)},$ (67b)
where we used a uniform distribution $\rho(\bm{x})=|\Omega|^{-1}=3/(4\pi(L^{3}-R^{3}))$. Then, one solves Eqs. (11) with $m=2$ for the second moment of FRTs:
$\langle\tau^{2}\rangle(r)$ $=\frac{1}{180D^{2}q^{2}rR^{4}}\bigg\{\begin{aligned} &40L^{6}(qR+1)\left(qrR-qR^{2}+r\right)-72L^{5}qR^{2}\left(qrR-qR^{2}+r\right)\\ &-20L^{3}R^{2}\left(q^{2}R\left(r^{3}-3r^{2}R+rR^{2}+R^{3}\right)+q\left(r^{3}-rR^{2}-2R^{3}\right)+4rR\right)\\ &+rR^{4}\left(q^{2}\left(3r^{4}-10r^{2}R^{2}+7R^{4}\right)+4q\left(5r^{2}R-7R^{3}\right)+40R^{2}\right)\bigg\},\end{aligned}$ (68a)
$\langle\tau^{2}\rangle(\circ)$ $=\frac{2}{315D^{2}q^{2}R^{4}\left(L^{3}-R^{3}\right)}\bigg\{\begin{aligned} &35L^{9}(qR+1)^{2}-126L^{8}qR^{2}(qR+1)+135L^{7}q^{2}R^{4}-105L^{6}R^{3}\\ &-63L^{5}qR^{5}(qR-2)+21L^{3}R^{6}(qR(qR-4)+5)+R^{9}(-2qR(qR-7)-35)\bigg\},\end{aligned}$ (68b)
thus determining the variances $\mathrm{Var}\{\tau\}(r)$ and $\mathrm{Var}\{\tau\}(\circ)$.
Moreover, by replacing $1/q$ by $\ell$ in $\langle\tau\rangle$, one has mean FCTs:
$\langle\mathcal{T}_{\ell}\rangle(r)$ $=\frac{(L^{3}-R^{3})\ell}{3DR^{2}}+\frac{2L^{3}(r-R)-r^{3}R+rR^{3}}{6DrR},$ (69a)
$\langle\mathcal{T}_{\ell}\rangle(\circ)$ $=\frac{(L^{3}-R^{3})\ell}{3DR^{2}}$
$+\frac{(L-R)^{2}\left(5L^{3}+6L^{2}R+3LR^{2}+R^{3}\right)}{15DR\left(L^{2}+LR+R^{2}\right)},$ (69b)
so that the cross-moments read
$\langle\tau\ell_{\tau}\rangle(r)=$ $\frac{2\left(L^{3}-R^{3}\right)}{3Dq^{2}R^{2}}+\frac{2L^{3}(r-R)-r^{3}R+rR^{3}}{6DqrR},$ (70a)
$\langle\tau\ell_{\tau}\rangle(\circ)=$ $\frac{2\left(L^{3}-R^{3}\right)}{3Dq^{2}R^{2}}$
$+\frac{(L-R)^{2}\left(5L^{3}+6L^{2}R+3LR^{2}+R^{3}\right)}{15DqR\left(L^{2}+LR+R^{2}\right)},$ (70b)
from which
$\langle\tau\ell_{\tau}\rangle-\langle\tau\rangle\langle\ell_{\tau}\rangle=\frac{|\Omega|}{|\partial\Omega|Dq^{2}}=\frac{L^{3}-R^{3}}{3Dq^{2}R^{2}}$ (71)
in both cases. Combining these results, one determines $C_{q}(\circ)$ from Eq. (15) for the shell approximation used in Fig. 7(b):
$C_{q}(\circ)$ $=\frac{5\sqrt{7}\left(L^{3}-R^{3}\right)^{2}}{\sqrt{\begin{aligned} &\bigg\{10\left(L^{3}-R^{3}\right)\big(35L^{9}(qR+1)^{2}-126L^{8}qR^{2}(qR+1)+135L^{7}q^{2}R^{4}-105L^{6}R^{3}\\ &-63L^{5}qR^{5}(qR-2)+21L^{3}R^{6}(qR(qR-4)+5)+R^{9}(-2qR(qR-7)-35)\big)\\ &-7\left(5L^{6}(qR+1)-9L^{5}qR^{2}+5L^{3}R^{3}(qR-2)+R^{6}(5-qR)\right)^{2}\bigg\}\end{aligned}}}\ .$ (72)
Appendix B Random packing of obstacles in two dimensions
As shown in Fig. 1(b), the area occupation ratio is 50%, including the reactive target and inert obstacles. The maximal radius is $1/(5\sqrt{2})\approx 0.141421$. The center of reactive target is located at $(0.0,0.0)$, while the coordinates of each obstacle’s center read (0.07068,0.3746)(-0.07068,0.3746), (0.1703,0.786)(0.1703,-0.786), (0.4497,0.7037)(0.4497,-0.7037), (0.343,0.4786)(-0.343,0.4786), (0.5462,0.02561)(-0.5462,-0.02561), $(0.3037,0.3557)$, (0.6671,0.5119)(-0.6671,-0.5119), (0.3806,0.6381)(-0.3806,-0.6381), (0.1114,0.4825)(-0.1114,-0.4825), $(0.4061,0.05633)$, (0.3541,0.2613)(0.3541,-0.2613), $(0.7377,0.01179)$, (0.6381,0.331)(-0.6381,0.331), (0.2481,0.2096)(-0.2481,-0.2096), $(0.4168,0.6535)$, (0.8433,0.1079)(-0.8433,0.1079), $(0.07206,0.7421)$, (0.7852,0.2465)(-0.7852,-0.2465), $(0.6535,0.3383)$, (0.2848,0.1658)(-0.2848,0.1658), (0.1314,0.8289)(-0.1314,-0.8289), (0.6499,0.3344)(0.6499,-0.3344), (0.2291,0.7764)(-0.2291,0.7764), (0.1874,0.4939)(0.1874,-0.4939).